# Rethinking Efficient and Effective Point-based Networks for Event Camera Classification and Regression: EventMamba

### v1

@@ Abstract
Event cameras, drawing inspiration from biological systems, efficiently detect changes in ambient light with low latency and high dynamic range while consuming minimal power. This innovative technology showcases the potential of bio-inspired design in developing advanced sensors for various applications. The most current approach to processing event data often involves converting it into frame-based representations, which is well-established in traditional vision. However, this approach neglects the sparsity of event data, loses fine-grained temporal information during the transformation process, and increases the computational burden, making it ineffective for characterizing event camera properties. In contrast, Point Cloud is a popular representation for 3D processing and is better suited to match the sparse and asynchronous nature of the event camera. Nevertheless, despite the theoretical compatibility of point-based methods with event cameras, the results show a performance gap that is not yet satisfactory compared to frame-based methods. In order to bridge the performance gap, we propose EventMamba, an efficient and effective Point Cloud framework that achieves competitive results even compared to the state-of-the-art (SOTA) frame-based method in both classification and regression tasks. This notable accomplishment is facilitated by our rethinking of the distinction between Event Cloud and Point Cloud, emphasizing effective temporal information extraction through optimized network structures. Specifically, EventMamba leverages temporal aggregation and State Space Model (SSM) based Mamba boasting enhanced temporal information extraction capabilities. Through a hierarchical structure, EventMamba is adept at abstracting local and global spatial features and implicit and explicit temporal features. By adhering to the lightweight design principle, EventMamba delivers impressive results with minimal computational resource utilization, demonstrating its efficiency and effectiveness. In the experiments, EventMamba exhibits SOTA point-based performance on four different scales of action recognition datasets and surpasses the majority of frame-based methods. Notably, EventMamba achieves, for the first time, a comparable accuracy to the SOTA results on the original RGB version of the HMDB51-DVS and UCF101-DVS datasets. Meanwhile, EventMamba distinguishes itself on both Camera Pose Relocalization (CPR) and eye-tracking regression tasks, outperforming all frame-based methods. EventMamba offers a paradigm for tackling event-based classification and regression tasks that can be derived to more tasks and real-world applications.
@@
Event Camera, Point Cloud Network, Mamba, Classification and Regression.
I Introduction
@@
Event cameras present an attractive option for deploying lightweight networks capable of real-time applications. Unlike traditional cameras, which sample light intensity at fixed intervals, event pixels independently respond to in-situ gradient light intensity changes, indicating the corresponding point’s movement in three-dimensional space [1]. This unique capability allows event cameras to capture sparse and asynchronous events that respond rapidly to changes in brightness. Event cameras demonstrate remarkable sensitivity to rapid light intensity changes, detecting log intensity in individual pixels and producing event outputs typically on the order of a million events per second (Meps). In comparison, standard cameras can only capture 60 frames per second, underscoring the exceptional speed of event cameras [2]. Furthermore, the distinctive operating principle of event cameras grants them superior power efficiency, low latency, less redundant information, and high dynamic range. Consequently, event cameras have found diverse applications in high-speed target tracking, simultaneous localization and mapping (SLAM) [3], as well as in industrial automation.
@@
Event Cloud is a data structure that encapsulates the coordinates, timestamps, and polarity of each event occurrence [4]. The temporal stream can discern fast behaviors with microsecond resolution [5]. Several studies have strived to augment and refine the problem of processing dynamic motion information from event cameras. Figure 2 (a) illustrates the frame-based method, which partitions the Event Cloud data into small chunks and accumulates the events for each time interval onto the image plane [6]. The grayscale on the plane denotes the intensity of the event occurrence at that point. However, the frame-based method disregards the fine-grained temporal information within a frame and may generate blurred images in fast-moving action scenarios. Moreover, this approach escalates the data density and processing volume, forfeiting the sparse data benefits of an event-based camera. Consequently, the frame-based method is unsuitable for scenarios with fast-moving actions and cannot be implemented with limited computational resources.
In order to overcome the issues discussed above, alternative approaches, such as the point-based method have been proposed [7, 8]. Point Cloud is a collection of 3D points $(x,y,z)$ that represents the shape and surface of an object or environment and is often used in lidar and depth cameras [9]. Event Cloud is a collection of events $(x,y,t,p)$ generated by event cameras, $t$ represents timestamps and $p$ is the polarity. By treating each event’s temporal information as the third dimension, event inputs $(x,y,t)$ can be transformed into points and aggregated into a pseudo-Point Cloud [7, 10, 11]. This method is well-suited for analyzing event data, which can then be directly fed into the Point Cloud network, as shown in Figure 2 (b). By doing so, the data conversion process is greatly simplified, and the fine-grained temporal information is preserved [7].
Nonetheless, contemporary point-based approaches still exhibit a discernible disparity in performance compared to frame-based approaches. Although the Event Cloud and the Point Cloud are identical in shape, the information they represent is inconsistent as mentioned above. Merely transplanting the most advanced Point Cloud network proves insufficient in extracting Event Cloud features, as the permutation invariance and translation invariance inherent in Point Cloud do not directly translate to the domain of Event Cloud [11]. Meanwhile, Event Cloud encapsulates a wealth of temporal information vital for event-based tasks. Consequently, addressing these distinctions necessitates the development of a tailored neural network module to effectively leverage the temporal features $t$ within the Event Cloud. This requests not only implicit temporal features, wherein $t$ and $x,y$ are treated as the same representation, but also explicit temporal features, where $t$ serves as the index of the features.
@@
Figure 1: The gain in three datasets of our work is visualized. EventMamba achieves SOTA in the point-based method and maintains very high efficiency. The shapes utilized to illustrate the results correspond to different input data formats, with varying colors denoting FLOPs.
@@
On the other hand, our primary goal is to democratize computer vision technology by making it accessible to a wider range of devices and applications. It is indeed accurate that the conventional frame-based method generally displays negligible processor power consumption. However, with the advent of DVS technology, the power consumption of sensors has been drastically reduced by 90%-95%, rendering it much lower than that of conventional ANN processors. Consequently, decreasing the algorithm’s power consumption has the potential to significantly reduce overall power consumption and enable the deployment of numerous battery-powered or resource-constrained devices, such as mobile phones, drones, and IoT devices. To achieve this goal, we introduce a novel Point Cloud framework, named EventMamba, for event-based action recognition [8], CPR, and eye tracking tasks, as shown in Figure 3. EventMamba is characterized by its lightweight parameterization and computational demands, making it ideal for resource-constrained devices. It features preprocessing, a hierarchy structure, and a classifier or regressor. First, the Event Cloud is divided into many sliding windows. The classification task is sampled through the subwindows [8], while the regression task is straightforward in accumulating enough points for processing [12]. Then, the sampled points are fed into a hierarchy structure to abstract high-level features. The hierarchy structure comprises local spatial, global spatial, and explicit temporal extractors. The feature extractors are interconnected through temporal aggregation. Finally, the general Multi-Layer Perceptron (MLP) is used for classification and regression. The main contributions of our work are:
@@
  • We directly process the raw data obtained from the event cameras, preserving the sparsity and fine-grained temporal information.
@@
  • We strictly arrange the events in chronological order, which can satisfy the prerequisites to extract the temporal features absent in Point Cloud.
@@
  • We employ temporal aggregation and Mamba modules to capture explicit temporal features during the hierarchy structure effectively.
@@
  • We propose a lightweight framework with a generalization ability to handle different-size event-based classification and regression tasks.
@@
II Related work
II-A Frame-based Method
Frame-based methods have been extensively employed for event data processing. These methods convert events into a density map by integrating the number of events within a specific time interval, which can be visualized as a 2D image [14]. The image can be directly fed into traditional feature extraction techniques as shown in Figure 2 for benchmarking purposes. The following will offer a comprehensive review of frame-based methods applied to two distinct tasks.
II-A1 Action recognition
The integration of the number of events within time intervals at each pixel to form a visual representation of a scene has been extensively studied, as documented in [14]. In cases where events repeatedly occur at the same spatial location, the resulting frame information is accumulated at the corresponding address, thus generating a grayscale representation of the temporal evolution of the scene. This concept is further explored in related literature, exemplified by the Event Spike Tensor (EST) [6] and the Event-LSTM framework [15], both of which demonstrate methods for converting sparse event data into a structured two-dimensional temporal-spatial representation, effectively creating a time-surface approach.
II-A2 Camera Pose Relocalization
PoseNet [19] introduced transfer learning by replacing the classifier of GoogLeNet [20] with a regressor specifically designed for absolute pose regression. Building upon this end-to-end concept, researchers have further improved the PoseNet framework by integrating advanced backbone architectures like VGG [21] and ResNet [22]. Alternative strategies have also been explored, such as incorporating LSTM networks as a substitute for the localizer or meticulously designing customized loss functions [23, 24, 25]. Additionally, the Auxiliary Learning methods have significantly contributed to boosting the network’s performance [26, 27, 28]. However, this end-to-end structure often suffers from overfitting, which can adversely impact performance.
Recently, hybrid approaches have emerged, combining learning-based techniques for local problems with traditional image retrieval or structure-based pipelines. NNnet [29] and RelocaNet [30] employed a backbone architecture to encode images and estimate poses by retrieving nearest neighbor images. On the other hand, approaches like DSAC series [31] took a different route by utilizing convolutional neural networks to regress scene coordinates. They further employed differentiable Random Sample Consensus (RANSAC) [32] for precise pose estimation, resulting in exceptional levels of accuracy. While these approaches did indeed yield a substantial reduction in estimation errors, they came at the cost of increased computational resources and higher inference latency. Conversely, the hybrid method shows promise but presents challenges due to its heavy reliance on extensive network support and demanding computational power to execute matching and retrieval algorithms. Such requirements pose significant obstacles, especially for edge devices.
II-A3 Eye-tracking
@@
In summary, frame-based methods have indeed showcased their prowess in event processing tasks, but they suffer from several limitations. For instance, the processed frame sizes are typically larger than those of the original Event Cloud data, resulting in high computational costs and large model sizes. Furthermore, the aggression process introduces significant latency [6], which hinders the use of event cameras in real-time human-machine interaction applications.
II-B Point-based Method
@@
III Methodology
III-A Event Cloud Preprocessing
@@
The symbol $R$ represents the time interval of the sliding window, where $j$ and $l$ denote the start and end event index of the sequence, respectively. $m$ represents the number of sliding windows into which the sequence of events $\mathcal{E}$ is divided. Before being fed into the neural network, $P_{i}$ also needs to undergo downsampling and normalization. Downsampling is to unify the number of Point Cloud, with the number of points in a sample defined as $N$, which is 1024 or 2048 in EventMamba. Additionally, we define the resolution of the event camera as $w$ and $h$. The normalization process is described by the following equation:
$PN_{i}=(\frac{X_{i}}{w},\frac{Y_{i}}{h},\frac{T_{i}-t_{j}}{t_{l}-t_{j}}),$ (3)
$X_{i},Y_{i},T_{i}=\{x_{j},\dots,x_{l}\},\{y_{j},\dots,y_{l}\},\{t_{j},\dots,t_{l}\},$ (4)
The $X,Y$ is divided by the resolution of the event camera. To normalize $T$, we subtract the smallest timestamp $t_{j}$ of the window and divide it by the time difference $t_{l}-t_{j}$, where $t_{l}$ represents the largest timestamp within the window. After pre-processing, Event Cloud is converted into the pseudo-Point Cloud, which comprises explicit spatial information $(x,y)$ and implicit temporal information $t$. It is worth mentioning that all events in $PN$ are strictly ordered according to the value of $t$, ensuring that the network can explicitly capture time information.
III-B Hierarchy Structure
@@
where $F$ is the extracted Point Cloud feature, which can be sent into the classifier or regressor to complete the different tasks. The model has the ability to extract global and local features and has a hierarchy structure [11]. And the dimension of $LocalFE$ input $\in$ is $[B,N,D]$, $m$ is the number of stages for hierarchy structure, $Agg$ is the operation of aggregation, $\mathcal{P}$ is the output of $LocalFE$, and the dimension of $GlobalFE$ input $\in$ $[B,N^{\prime},D^{\prime}]$, as illustrated in Figure 3. $B$ is the batch size, $D$ is the dimension of features, and $N$ is the number of $Centroid$ in the overall Point Cloud during the first stage of the hierarchy structure.
From a macro perspective, $LocalFE$ is mainly used to extract the local features of the spatial domain and pay attention to local geometric features through the farthest point sampling and grouping method. Temporal aggregation is responsible for aggregating local features on the timeline. $GlobalFE$ mainly synthesizes the extracted local features and captures the explicit temporal relationship between different groups. Details of the three modules are described below.
III-B1 $LocalFE$ Extractor

### v2

@@ Abstract
Event cameras, drawing inspiration from biological systems, efficiently detect changes in ambient light with low latency and high dynamic range while consuming minimal power. The most current approach to processing Event Cloud often involves converting it into frame-based representations, which neglects the sparsity of events, loses fine-grained temporal information, and increases the computational burden. In contrast, Point Cloud is a popular representation for 3D processing and serves as an alternative method to address the limitations mentioned above. Nevertheless, previous point-based methods show an unsatisfactory performance compared to the frame-based method. In order to bridge the gap, we propose EventMamba, an efficient and effective framework based on Point Cloud representation that achieves competitive results even compared to the state-of-the-art (SOTA) frame-based method in both classification and regression tasks. This notable accomplishment is facilitated by our rethinking of the distinction between Event Cloud and Point Cloud, emphasizing effective temporal information extraction through optimized network structures. Specifically, EventMamba leverages temporal aggregation and State Space Model (SSM) based Mamba boasting enhanced temporal information extraction capabilities. Through a hierarchical structure, EventMamba is adept at abstracting local and global spatial features and implicit and explicit temporal features. By adhering to the lightweight design principle, EventMamba delivers impressive results with minimal computational resource utilization, demonstrating its efficiency and effectiveness. In the experiments, EventMamba exhibits SOTA point-based performance on five different scales of action recognition datasets and distinguishes itself on both Camera Pose Relocalization (CPR) and eye-tracking regression tasks, outperforming all frame-based methods.
@@
Event Camera, Point Cloud Network, Spatio-temporal, Mamba, Classification and Regression.
1 Introduction
@@
Event cameras present an attractive option for deploying lightweight networks capable of real-time applications. Unlike traditional cameras, which sample light intensity at fixed intervals, event pixels independently respond to in-situ gradient light intensity changes, indicating the corresponding point’s movement in three-dimensional space [1]. This unique capability allows event cameras to capture sparse and asynchronous events that respond rapidly to changes in brightness. Event cameras demonstrate remarkable sensitivity to rapid light intensity changes, detecting log intensity in individual pixels and producing event outputs typically on the order of a million events per second (eps). In comparison, standard cameras can only capture 60 frames per second, underscoring the exceptional speed of event cameras [2]. Furthermore, the distinctive operating principle of event cameras grants them superior power efficiency, low latency, less redundant information, and high dynamic range. Consequently, event cameras have found diverse applications in high-speed target tracking, simultaneous localization and mapping (SLAM), as well as in industrial automation.
@@
Event Cloud is a data structure that encapsulates the coordinates, timestamps, and polarity of each event occurrence [3]. The temporal stream can discern fast behaviors with microsecond resolution [4]. Several studies have strived to augment and refine the problem of processing dynamic motion information from event cameras. Figure 2 (a) illustrates the frame-based method, which partitions the Event Cloud into small chunks and accumulates the events for each time interval onto the image plane [5]. The grayscale on the plane denotes the intensity of the event occurrence at that coordinate. However, the frame-based method disregards the fine-grained temporal information within a frame and may generate blurred images in fast-moving action scenarios. Moreover, this approach escalates the data density and computes even in event-free regions, forfeiting the sparse data benefits of an event-based camera. Consequently, the frame-based method is unsuitable for scenarios with fast-moving actions and cannot be implemented with limited computational resources.
@@
Figure 1: The gain in three datasets is visualized. EventMamba achieves SOTA in the point-based method and maintains very high efficiency. The shapes illustrate the results corresponding to different data formats, with varying colors denoting Floating Point Operations Per Second (FLOPs).
In order to overcome the issues discussed above, alternative approaches, such as the point-based method have been proposed [6, 7, 8]. Point Cloud is a collection of 3D points $(x,y,z)$ that represents the shape and surface of an object or environment and is often used in lidar and depth cameras [9]. Event Cloud is a collection of events $(x,y,t,p)$ generated by event cameras, $t$ represents timestamps and $p$ is the polarity. By treating each event’s temporal information as the third dimension, event inputs $(x,y,t)$ can be regarded as points and aggregated into a pseudo-Point Cloud [6, 10, 11]. This method is well-suited for analyzing Event Cloud, which can then be directly fed into the Point Cloud network, as shown in Figure 2 (b). By doing so, the data format transformation process is greatly simplified, and the fine-grained temporal information and sparse property are preserved [6].
Nonetheless, contemporary point-based approaches still exhibit a discernible disparity in performance compared to frame-based approaches. Although the Event Cloud and the Point Cloud are identical in shape, the information they represent is inconsistent, as mentioned above. Merely transplanting the most advanced Point Cloud network proves insufficient in extracting Event Cloud features, as the permutation invariance and transformation invariance inherent in Point Cloud do not directly translate to the domain of Event Cloud [11]. Meanwhile, Event Cloud encapsulates a wealth of temporal information vital for event-based tasks. Consequently, addressing these distinctions necessitates the development of a tailored neural network module to effectively leverage the temporal features $t$ within the Event Cloud. This requests not only implicit temporal features, wherein $t$ and $x,y$ are treated as the same representation, but also explicit temporal features, where $t$ serves as the index of the features.
@@
On the other hand, our primary goal is to democratize computer vision technology by making it accessible to a wider range of devices and applications. It is indeed accurate that the conventional frame-based method generally displays negligible processor power consumption. However, with the advent of DVS technology, the power consumption of sensors has been drastically reduced by 90%-95% [12], rendering it much lower than that of conventional ANN processors. As a result, decreasing the algorithm’s power consumption has the potential to significantly reduce overall power consumption and enable the deployment of numerous battery-powered or resource-constrained devices, such as mobile phones, drones, and IoT devices. To achieve this goal, we introduce a novel Point Cloud framework, named EventMamba, for event-based action recognition [7], CPR, and eye-tracking tasks, whose gains are shown in Figure 1. EventMamba is characterized by its lightweight parameterization and computational demands, making it ideal for resource-constrained devices. It features preprocessing, a hierarchy structure, and a classifier or regressor. First, the Event Cloud is divided into many sliding windows. The classification task is sampled through the subwindows [7], while the regression task is straightforward in accumulating enough points for processing [8]. Then, the sampled points are fed into a hierarchy structure to abstract high-level features. The hierarchy structure comprises local spatial, global spatial, and explicit temporal extractors. The feature extractors are interconnected through temporal aggregation. Finally, the general Multi-Layer Perceptron (MLP) is used for classification and regression. The main contributions of our work are:
@@
1. We directly process the raw data obtained from the event cameras, preserving the sparsity and fine-grained temporal information.
@@
2. We strictly arrange the events in chronological order, which can satisfy the prerequisites to extract the temporal features absent in vanilla Point Cloud Network.
@@
3. We employ temporal aggregation and Mamba modules to capture explicit temporal features during the hierarchy structure effectively.
@@
4. We propose a lightweight framework with a generalization ability to handle different-size event-based classification and regression tasks.
2 Related Works
2.1 Frame-based Method
Frame-based methods have been extensively employed for event data processing. These methods convert events into a density map by integrating the number of events within a specific time interval, which can be visualized as a 2D image [13]. The image can be directly fed into traditional feature extraction techniques as shown in Figure 2 (a) for benchmarking purposes. The following will offer a comprehensive review of frame-based methods applied to three distinct tasks.
2.1.1 Action Recognition
Action recognition is a critical task with diverse applications in anomaly detection, entertainment, and security. Most event-based approaches often have borrowed from traditional camera-based approaches. The integration of the number of events within time intervals at each pixel to form a visual representation of a scene has been extensively studied, as documented in [13]. In cases where events repeatedly occur at the same spatial location, the resulting frame information is accumulated at the corresponding address, thus generating a grayscale representation of the temporal evolution of the scene. This concept is further explored in related literature, exemplified by the Event Spike Tensor (EST) [5] and the Event-LSTM framework [14], both of which demonstrate methods for converting sparse event data into a structured two-dimensional temporal-spatial representation, effectively creating a time-surface approach.
@@ 2.1.1 Action Recognition
2.1.2 Camera Pose Relocalization
CPR task is an emerging application in power-constrained devices and has gained significant attention. It aims to train several scene-specific neural networks to accurately relocalize the camera pose within the original scene used for training. RGB-based CPR networks have been extensively studied [19]. PoseNet [20] introduced transfer learning by replacing the classifier of GoogLeNet [21] with a regressor specifically designed for absolute pose regression. Building upon this end-to-end concept, researchers have further improved the PoseNet framework by integrating advanced backbone architectures like VGG [22] and ResNet [23]. Alternative strategies have also been explored, such as incorporating LSTM networks as a substitute for the localizer or meticulously designing customized loss functions [24, 25, 26]. Additionally, the Auxiliary Learning methods have significantly contributed to boosting the network’s performance [27, 28, 29]. However, this end-to-end structure often suffers from overfitting, which can adversely impact performance. Recently, hybrid approaches have emerged, combining learning-based techniques for local problems with traditional image retrieval or structure-based pipelines. DSAC series, for instance, achieve high pose estimation accuracy [30, 31], but come with increased computational costs and latency, especially for edge devices.
Event-based CPR methods often derive from the frame-based CPR network. SP-LSTM [32] employed the stacked spatial LSTM networks to process event images, facilitating a real-time pose estimator. To address the inherent noise in event images, Jin Yu et al.[33] proposed a network structure combining denoise networks, convolutional neural networks, and LSTM, achieving good performance under complex working conditions. In contrast to the aforementioned methods, a novel representation named Reversed Window Entropy Image (RWEI) [34] is introduced, which is based on the widely used event surface [35] and serves as the input to an attention-based DSAC* pipeline [30] to achieve SOTA results. However, the computationally demanding architecture involving representation transformation and hybrid pipeline poses challenges for real-time execution. Additionally, all existing methods ignore the fine-grained temporal feature of the event cameras, and accumulate events into frames for processing, resulting in unsatisfactory performance.
2.1.3 Eye-tracking
@@
In summary, frame-based methods have indeed showcased their prowess in event processing tasks, but they suffer from several limitations. For instance, the processed frame sizes are typically larger than those of the original Event Cloud, resulting in high computational costs and large model sizes. Furthermore, the aggression process introduces significant latency [5], which hinders the use of event cameras in real-time human-machine interaction applications. Despite substantial endeavors aimed at mitigating these challenges [41], changing the representation stands as the suitable approach to delving deep into the core of the issue.
2.2 Point-based Method
@@
3 Methodology
3.1 Event Cloud Preprocessing
@@
The symbol $R$ represents the time interval of the sliding window, where $j$ and $l$ denote the start and end event index of the sequence, respectively. $L$ represents the number of sliding windows into which the sequence of events $\mathcal{E}$ is divided. Before being fed into the neural network, $P_{i}$ also needs to undergo downsampling and normalization. Downsampling is to unify the number of Point Cloud, with the number of points in a sample defined as $N$, which is 1024 or 2048 in EventMamba. Additionally, we define the resolution of the event camera as $w$ and $h$. The normalization process is described by the following equation:
$PN_{i}=(\frac{X_{i}}{w},\frac{Y_{i}}{h},\frac{TS_{i}-t_{j}}{t_{l}-t_{j}}),$ (3)
$X_{i},Y_{i},TS_{i}=\{x_{j},\dots,x_{l}\},\{y_{j},\dots,y_{l}\},\{t_{j},\dots,t_{l}\},$ (4)
The $X,Y$ is divided by the resolution of the event camera. To normalize timestamps $TS$, we subtract the smallest timestamp $t_{j}$ of the window and divide it by the time difference $t_{l}-t_{j}$, where $t_{l}$ represents the largest timestamp within the window. After pre-processing, Event Cloud is converted into the pseudo-Point Cloud, which comprises explicit spatial information $(x,y)$ and implicit temporal information $t$. It is worth mentioning that all events in $PN$ are strictly ordered according to the value of $t$, ensuring that the network can explicitly capture time information.
3.2 Hierarchy Structure
@@
where $F$ is the extracted spatio-temporal Event Cloud feature, which can be sent into the classifier or regressor to complete the different tasks. The model has the ability to extract global and local features and has a hierarchy structure [11]. And the dimension of $LocalFE$ input $\in$ is $[B,T_{i-1},D_{i-1}]$, $m$ is the number of stages for hierarchy structure, $TAG$ is the operation of temporal aggregation, $\mathcal{S}$ is the output of $LocalFE$, and the dimension of $GlobalFE$ input $\in$ $[B,T_{i},D_{i}]$, as illustrated in Figure 3. $B$ is the batch size, $D$ is the dimension of features, and $T$ is the number of $Centroid$ in the overall Point Cloud during the current stage of the hierarchy structure.
From a macro perspective, $LocalFE$ is mainly used to extract the local features of the spatial domain and pay attention to local geometric features through the farthest point sampling and grouping method. Temporal aggregation ($TAG$) is responsible for aggregating local features on the timeline. $GlobalFE$ mainly synthesizes the extracted local features and captures the explicit temporal relationship between different groups. Details of the three modules are described below.
3.2.1 $LocalFE$ Extractor